A rectangle is cut by 6 lines parallel and 8 lines perpendicular to the base. Find the total number of rectangles thus formed
Answer by ikleyn(52776)

.
A rectangle is cut by 6 lines parallel and 8 lines perpendicular to the base.
Find the total number of rectangles thus formed
~~~~~~~~~~~~~~~~

First 6 cuts made 7 = 6+1  strips parallel to the base.

Next 8 cuts perpendicular to the base cut each of 7 parallel strips in 9 = 8+1 small rectangles.


Total number of small cut rectangles is the product 7*9 = 63.    ANSWER